OpenStudy (anonymous):

Solve the equation. Show your work. 1/a=6/18

OpenStudy (pfenn1):

To solve the equation you want to rearrange the equation to get just a on one side of the equal sign and the rest of the numbers on the other side. Do you know how to do this?

OpenStudy (anonymous):

\[\frac{1}{a}=\frac{6}{18}=\frac{1}{3}\] now you solve with your eyeballs

OpenStudy (anonymous):

@saifoo.khan how did you find that????

OpenStudy (saifoo.khan):

@satellite73 : shh. ;) i'll inbox.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

...

OpenStudy (anonymous):

no